Solve the inequality √(4x2-8x+7) ≤ -2x



1
Expert's answer
2021-10-25T16:19:32-0400

"\sqrt{4x^2-8x+7}\leq-2x"


"\sqrt{4x^2-8x+7} ≤ -2x,
 x\isin\reals"


Separate into two possible case


"\sqrt{4x^2-8x+7}≤ -2x,-2x≥0"


"\sqrt{4x^2-8x+7}≤ -2x,-2x<0"


Solve the inequality for x


"x≥\frac{7}{8}, -2x≥0"


"\sqrt{4x^2-8x+7}≤ -2x,-2x<0"


Since the left-hand side is always positive or zero, and the right-hand side is always negative, the statement is false for any value of x


"x≥\frac{7}{8},x≤0"


"\sqrt{4x^2-8x+7}≤ -2x,-2x<0"



"x\isin\varnothing" , "-2x<0"


"x\isin\varnothing,x>0"


Find the intersection


"x\isin\varnothing"

"x\isin\varnothing,x>0"


The union

"x\isin\varnothing"

That is, there is no solution.